# Variational Continual Test-Time Adaptation

### v1

@@ 5.1 Experimental setting
Dataset. In our experiments, we employ the CIFAR10C, CIFAR100C, and ImageNetC datasets as benchmarks to assess the robustness of classification models. Each dataset comprises 15 distinct types of corruption, each applied at five different levels of severity (from 1 to 5). These corruptions are systematically applied to test images from the original CIFAR10 and CIFAR100 datasets, as well as validation images from the original ImageNet dataset. For simplicity in tables, we use C1 to C15 to represent the 15 types of corruption, i.e., C1: Gaussian, C2: Shot, C3: Impulse C4: Defocus, C5: Glass, C6: Motion, C7: Zoom, C8: Snow, C9: Frost, C10: Fog, C11: Brightness, C12: Contrast, C13: Elastic, C14: Pixelate, C15: Jpeg.
@@
Pretrained Model. Following previous studies (Wang et al., 2020, 2022), we adopt pretrained WideResNet-28 (Zagoruyko & Komodakis, 2016) model for CIFAR10to-CIFAR10C, pre-trained ResNeXt-29 (Xie et al., 2017) for CIFAR100-to-CIFAR100C, and standard pre-trained ResNet-50 (He et al., 2016) for ImageNet-to-ImagenetC. Note in our VCoTTA (Wang et al., 2022), we further warm up the pretrained model to obtain the stochastic dynamics for each dataset. Similar to CoTTA, we update all the trainable parameters in all experiments. The augmentation number is set to 32 for all compared methods using the augmentation strategy.
Table 4: Ablation study on Variational Warm-Up (VWU) and Symmetric Cross-Entropy (SCE) under severity 5.
No. VWU SCE CIFAR10C CIFAR100C ImageNetC
1 18.4 31.5 68.1
2 $\surd$ 17.1 31.2 68.3
3 $\surd$ 13.9 28.8 64.2
4 $\surd$ $\surd$ 13.1 28.4 64.7
5.2 Methods to be compared
We compare our VCoTTA with multiple state-of-the-art (SOTA) methods. Source denotes the baseline pretrained model without any adaptation. BN (Li & Hoiem, 2017; Schneider et al., 2020) keeps the network parameters frozen, but only updates Batch Normalization. TENT (Wang et al., 2020) updates via Shannon entropy for unlabeled test data. CoTTA (Wang et al., 2022) builds the MT structure and uses randomly restoring parameters to the source model. SATA (Chakrabarty et al., 2023) modifies the batch-norm affine parameters using source anchoring-based self-distillation to ensure the model incorporates knowledge of newly encountered domains while avoiding catastrophic forgetting. SWA (Yang et al., 2023) refines the pseudo-label learning process from the perspective of the instantaneous and long-term impact of noisy pseudo-labels. PETAL (Brahma & Rai, 2023) tries to estimate the uncertainty in CTTA, which is similar to BNN, but it ignores the prior shift problem. All compared methods adopt the same backbone, pretrained model and hyperparameters.
5.3 Comparison results
We show the major comparisons with the SOTA methods in Tables 1, 2 and 3. We have the following observations. First, no adaptation at the test time (Source) suffers from serious domain shift, which shows the necessity of the CTTA. Second, traditional TTA methods that ignore the continual shift in test time perform poorly such as TENT and BN. We also find that simple Shannon entropy is effective in the first several domain shifts, especially in complex 1,000-classes ImageNetC, but shows significant performance drops in the following shifts. Third, the mean-teacher structure is very useful in CTTA, such as CoTTA and PETAL, which means that the pseudo-label is useful in domain shift. In the previous method, the error accumulation leads to the unreliable pseudo labels, then the model may get more negative transfers in CTTA along the timeline. The proposed VCoTTA outperforms other methods on all the three datasets, such as 13.1% vs. 15.3% (SWA) on CIFAR10C, 28.4% vs. 30.0% (SWA) on CIFAR100C and 64.2% vs. 66.7% (CoTTA) on ImageNetC. We hold the opinion that the prior will inevitably drift in CTTA, but VCoTTA slows down the process via the prior mixture. We also find that the superiority is more obvious in the early adaptation, which may be influenced by the different corruption orders. We analyze the order problem in Sec. 5.8.
5.4 Ablation study
We evaluate the two components in Table 4, i.e., the Variational Warm-Up (VWU) and the Symmetric Cross-Entropy (SCE) via ablation. The ablation results show that the two components are both important for VCoTTA. First, the VWU is used to inject stochastic dynamics into an off-the-shelf pretrained model. Without the VWU, the performance of VCoTTA drops to 18.4% from 13.9% on CIFAR10C, 31.5% from 28.8% on CIFAR100C and 68.1% from 64.2% on ImageNetC. Also, the SCE can further improve the performance on CIFAR10C and CIFAR100C, because SCE balances the gradient for high and low confidence predictions. We also find that SCE is ineffective for complex ImageNetC, and the reason may be the class sensitivity imbalance, causing the model to lean more towards one direction during optimization.
5.5 Mixture of priors
In Sec. 4.2, we introduce a Gaussian mixture strategy, where the current prior is approximated as the weighted sum of the source prior and the teacher prior. The weights are determined by computing the entropy over multiple augmentations of two models. To assess the effectiveness of these weights, we compare them with three naive weighting configurations: using only the source model, using only the teacher model, and a simple average with equal weights for both models. The results, as presented in Table 5, reveal that relying solely on the source model or the teacher model (i.e., weighting with $(1,0)$ and $(0,1)$) results in suboptimal performance. Additionally, naive weighting with equal contributions from both models (i.e., $(0.5,0.5)$) proves ineffective for CTTA due to the inherent uncertainty in both models. In contrast, the proposed adaptive weights for the Gaussian mixture in CTTA demonstrate its effectiveness. This underscores the significance of striking a balance between the two prior models in an unsupervised environment. The trade-off implies the need to discern when the source model’s knowledge is more applicable and when the teacher model’s shifting knowledge takes precedence.
Table 5: Different weights for mixture of priors (Severity 5).
No. $p_{1}(\boldsymbol{\theta})$ $\bar{p}_{t}(\boldsymbol{\theta})$ CIFAR10C CIFAR100C ImageNetC
@@ 5.5 Mixture of priors
3 0.5 0.5 14.7 31.3 67.0
4 $\alpha$ $1-\alpha$ 13.1 28.4 64.7
5.6 Uncertainty estimation
To evaluate the uncertainty estimation, we use negative loglikelihood (NLL) and Brier Score (BS) (Brier, 1950). Both NLL and BS are proper scoring rules (Gneiting & Raftery, 2007), and they are minimized if and only if the predicted distribution becomes identical to the actual distribution:
@@ 5.6 Uncertainty estimation
where $\mathcal{D}^{\text{test}}$ denotes the test set, i.e., the unsupervised test dataset $\mathcal{U}$ with labels. We evaluate NLL and BS with a severity level of 5 for all corruption types, and the compared results with SOTAs are shown in Table 6. We have the following observations. First, most methods suffer from low confidence in terms of NLL and BS because of the drift priors, where the model is unreliable gradually, and the error accumulation makes the model perform poorly. Our approach outperforms most other approaches in terms of NLL and BS, demonstrating the superiority in improving uncertainty estimation. We also find that PETAL (Brahma & Rai, 2023) shows good NLL and BS, because PETAL forces the prediction over-confident to unreliable priors, thus PETAL shows unsatisfactory results on adaptation accuracy, such as 31.5% vs. 28.4% (Ours) on CIFAR100C.
Table 6: Uncertainty estimation via NLL and BS.
@@
5.7 Gradually corruption
Following CoTTA (Wang et al., 2022), we show gradual corruption results instead of constant severity in the major comparison, and the results are reported in Table 7. Specifically, each corruption adopts the gradual changing sequence: $1\rightarrow 2\rightarrow 3\rightarrow 4\rightarrow 5\rightarrow 4\rightarrow 3\rightarrow 2\rightarrow 1$, where the severity level is the lowest 1 when corruption type changes, therefore, the type change is gradual. The distribution shift within each type is also gradual. Under this situation, our VCoTTA also outperforms other methods, such as 8.9% vs. 10.5% (PETAL) on CIFAR10C, and 24.4% vs. 26.3% (CoTTA) on CIFAR100C. The results show that the proposed VCoTTA based on BNN is also effective when the distribution change is uncertain.
Table 7: Gradually changing on severity.
No. CIFAR10C CIFAR100C ImageNetC
@@ 5.7 Gradually corruption
5.8 Different orders of corruption
As we discuss in the major comparisons (see Sec 5.3), the performance may be affected by the corruption order. To provide a more comprehensive evaluation of the matter of the order, we conduct 10 different orders from Sec 5.3, and show the average performance of all compared methods. 10 independent random orders of corruption are all under the severity level of 5. The results are shown in Table 8. We find that the order of corruption is minor on simple datasets such as CIFAR10C and CIFAR100C, but small std on difficult datasets such as ImageNetC. The proposed VCoTTA outperforms other methods on the average error of CIFAR10C and CIFAR100C under 10 different corruption orders, which shows the effectiveness of the prior calibration in CTTA. Moreover, VCoTTA has comparable results with PETAL on ImageNetC, but smaller std over 10 orders, which shows the robustness of the proposed method.
Table 8: Comparisons over 10 orders (avg $\pm$ std).
Method CIFAR10C CIFAR100C ImageNetC
CoTTA 17.3$\pm$0.3 32.2$\pm$0.3 63.4$\pm$3.0
PETAL 16.0$\pm$0.1 33.8$\pm$0.3 62.7$\pm$2.6
VCoTTA 13.1$\pm$0.1 28.2$\pm$0.2 62.8$\pm$1.1
5.9 Corruption loops
Figure 3: 10 loops under a same corruption order (CIFAR10C).
5.10 Hyperparameter analysis
In this subsection, we analyze the two major hyperparameters in VCoTTA, i.e., the margin $\epsilon$ in Eq. (13) and the number of augmentation in Eq. (10). We experimentally validate these hyperparameters with more choices. Experimental results are shown in Table 9 and Table 10. The results indicate that different datasets may require different margins as thresholds to control confidence. Additionally, increasing the number of augmentations can enhance effectiveness, but this hyperparameter ceases to have a significant impact after reaching 32.
Table 9: Analysis on confidence margin.

### v2

@@ VI-A1 Dataset
We use CIFAR10-to-CIFAR10C, CIFAR100-to-CIFAR100C, and ImageNet-to-ImageNetC as benchmarks to assess the robustness of classification models. Each dataset comprises 15 distinct types of corruption, each applied at severity 5. These corruptions are systematically applied to test images from the original CIFAR10 and CIFAR100 datasets, as well as validation images from the original ImageNet dataset. For simplicity in tables, we use C1 to C15 to represent the 15 types of corruption, i.e., C1: Gaussian, C2: Shot, C3: Impulse C4: Defocus, C5: Glass, C6: Motion, C7: Zoom, C8: Snow, C9: Frost, C10: Fog, C11: Brightness, C12: Contrast, C13: Elastic, C14: Pixelate, C15: Jpeg.
TABLE I: Classification error rates (%) for the standard CIFAR10-to-CIFAR10C and CIFAR100-to-CIFAR100C CTTA tasks. All models are evaluated under the highest corruption severity level (level 5) in an online adaptation setting. C1 to C15 denote the 15 corruption types described in Section VI-A. The best result is highlighted in bold, and the second-best result is underlined.
Method Venue C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 C11 C12 C13 C14 C15 Avg
CIFAR10-to-CIFAR10C
Source - 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.5 30.3 43.5
TENT [wang2020tent] ICLR 2021 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
Ada [AdaContrast] CVPR 2022 29.1 22.5 30.0 14.0 32.7 14.1 12.0 16.6 14.9 14.4 8.1 10.0 21.9 17.7 20.0 18.5
CoTTA [wang2022continual] CVPR 2022 24.3 21.3 26.6 11.6 27.6 12.2 10.3 14.8 14.1 12.4 7.5 10.6 18.3 13.4 17.3 16.2
EATA [EATA] ICML 2022 27.3 20.3 28.6 13.5 27.8 14.6 13.5 17.9 19.0 16.4 10.7 14.6 20.8 16.4 19.1 18.7
RoTTA [yuan2023robust] CVPR 2023 30.3 25.4 34.6 18.3 34.0 14.7 11.0 16.4 14.6 14.0 8.0 12.4 20.3 16.8 19.4 19.3
PETAL [brahma2023probabilistic] CVPR 2023 26.0 18.9 27.2 12.1 26.4 13.2 12.1 16.5 18.5 15.0 10.2 13.2 19.4 15.0 17.6 17.3
RMT [dobler2023robust] CVPR 2023 24.0 20.4 25.6 12.6 25.4 14.2 12.2 15.4 15.1 14.1 10.3 13.7 17.1 13.5 16.0 16.7
BeCOTTA [becotta] ICLR 2024 22.9 19.1 26.9 10.2 27.5 12.7 10.4 14.7 14.3 12.4 7.2 9.4 20.9 15.2 20.2 16.3
SANTA [chakrabarty2023sata] TMLR 2023 23.9 20.1 28.0 11.6 27.4 12.6 10.2 14.1 13.2 12.2 7.4 10.3 19.1 13.3 18.5 16.1
SWA [yang2023exploring] IJCAI 2023 23.9 20.5 24.5 11.2 26.3 11.8 10.1 14.0 12.7 11.5 7.6 9.5 17.6 12.0 15.8 15.3
DSS [wang2024continual] WACV 2024 24.1 21.3 25.4 11.7 26.9 12.2 10.5 14.5 14.1 12.5 7.8 10.8 18.0 13.1 17.3 16.0
OBAO [obao] ECCV 2024 23.6 19.9 26.0 11.8 25.3 13.2 10.9 14.3 13.5 12.7 9.0 11.9 17.1 12.7 15.9 15.8
PALM [palm] AAAI 2025 25.9 18.1 22.7 12.4 25.3 13.2 10.8 13.5 13.2 12.2 8.5 11.9 17.9 12.0 15.5 15.5
SURGEON [surgeon] CVPR 2025 25.3 20.0 27.4 12.0 28.0 13.7 11.1 15.3 14.3 14.1 8.2 11.8 19.2 13.5 18.7 16.8
TCA [CTTA-TCA] CVPR 2025 22.6 19.4 23.2 11.0 23.4 12.0 10.1 13.3 13.5 11.9 7.9 11.0 16.7 12.3 15.7 14.9
VCoTTA Ours 18.1 14.9 22.0 9.7 22.6 11.0 9.5 11.4 10.6 10.5 6.5 9.4 15.6 11.0 14.5 13.1
CIFAR100-to-CIFAR100C
Source - 73.0 68.0 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
TENT [wang2020tent] ICLR 2021 37.2 35.8 41.7 37.9 51.2 48.3 48.5 58.4 63.7 71.1 70.4 82.3 88.0 88.5 90.4 60.9
Ada [AdaContrast] CVPR 2022 42.3 36.8 38.6 27.7 40.1 29.1 27.5 32.9 30.7 38.2 25.9 28.3 33.9 33.3 36.2 33.4
CoTTA [wang2022continual] CVPR 2022 40.1 37.7 39.7 26.9 38.0 27.9 26.4 32.8 31.8 40.3 24.7 26.9 32.5 28.3 33.5 32.5
EATA [EATA] ICML 2022 46.7 34.5 42.4 25.9 36.8 31.6 25.3 35.6 33.8 41.1 22.6 28.6 36.6 28.4 33.5 33.6
RoTTA [yuan2023robust] CVPR 2023 49.1 44.9 45.5 30.2 42.7 29.5 26.1 32.2 30.7 37.5 24.7 26.9 32.5 28.3 33.5 32.5
PETAL [brahma2023probabilistic] CVPR 2023 42.1 38.6 38.7 28.1 38.1 28.1 27.5 31.2 30.7 37.2 26.4 28.7 29.4 28.6 31.6 32.8
RMT [dobler2023robust] CVPR 2023 40.5 36.1 36.3 27.7 33.9 28.5 26.4 29.0 29.0 32.5 25.1 27.4 28.2 26.3 29.3 30.4
BeCoTTA [becotta] ICLR 2024 42.1 38.0 42.2 30.2 42.9 31.7 29.8 35.1 33.9 38.5 27.9 32.0 36.7 31.6 39.9 35.5
SANTA [chakrabarty2023sata] TMLR 2023 36.5 33.1 35.1 25.9 34.9 27.7 25.4 29.5 29.9 33.1 23.6 26.7 31.9 27.5 35.2 30.3
SWA [yang2023exploring] IJCAI 2023 39.4 36.4 37.4 25.0 36.0 26.6 25.0 29.1 28.4 35.0 23.5 25.1 28.5 25.8 29.6 30.0
DSS [wang2024continual] WACV 2024 39.7 36.0 37.2 26.3 35.6 27.5 25.1 31.4 30.0 37.8 24.2 26.0 30.0 26.3 31.3 30.9
OBAO [obao] ECCV 2024 38.8 35.0 35.4 26.7 33.2 27.4 25.4 29.4 28.8 32.8 24.1 26.1 30.9 26.9 32.0 30.2
PALM [palm] AAAI 2025 37.3 32.5 34.9 26.2 35.3 27.5 24.6 29.1 29.2 34.1 23.5 27.0 31.1 26.6 34.1 30.2
SURGRON [surgeon] CVPR 2025 40.3 35.6 38.1 26.4 36.7 28.1 25.5 31.2 30.0 35.6 24.8 27.2 31.3 28.2 36.1 31.7
TCA [CTTA-TCA] CVPR 2025 38.5 36.0 36.6 25.8 34.8 27.2 25.1 30.5 27.8 30.5 24.1 25.7 27.3 26.6 30.5 29.8
VCoTTA Ours 35.3 32.8 38.9 23.8 34.6 25.5 23.2 27.5 26.7 30.4 22.1 23.0 28.1 24.2 30.4 28.4
VI-A2 Pretrained Model
@@
Following previous studies [wang2020tent, wang2022continual], we adopt pretrained WideResNet-28 [zagoruyko2016wide] model for CIFAR10to-CIFAR10C, pretrained ResNeXt-29 [xie2017aggregated] for CIFAR100-to-CIFAR100C, and standard pretrained ResNet-50 [he2016deep] for ImageNet-to-ImagenetC. Note in our VCoTTA [wang2022continual], we further warm up the pretrained model to obtain the stochastic dynamics for each dataset. Similar to CoTTA, we update all the trainable parameters in all experiments. The augmentation number is set to 32 for all compared methods that use the augmentation strategy.
VI-B Comparison Results
We compare our VCoTTA with multiple state-of-the-art (SOTA) methods, including TENT [wang2020tent], Ada [AdaContrast] CoTTA [CTTA-CoTTA], EATA [EATA], RoTTA [yuan2023robust] PETAL [brahma2023probabilistic], RMT [dobler2023robust], BeCOTTA [becotta], SANTA [chakrabarty2023sata], SWA [yang2023exploring], DSS [wang2024continual], OBAO [obao], PALM [palm], SURGEON [surgeon], and TCA [CTTA-TCA]. Source denotes the baseline pre-trained model without any adaptation. All compared methods adopt the same backbone, pre-trained model and hyperparameters.
We present major comparisons with state-of-the-art methods on CIFAR10C, CIFAR100C and ImageNetC, and summarize the main findings. First, source-only models without adaptation suffer heavily from domain shifts, which clearly demonstrates the necessity of CTTA. Second, traditional TTA methods that ignore continual shifts, such as TENT, perform poorly in this setting. Third, mean-teacher based approaches including CoTTA and PETAL confirm the usefulness of pseudo-labeling, although they remain vulnerable to error accumulation, which gradually produces unreliable pseudo labels and negative transfer as adaptation progresses. As shown in Table I, on CIFAR10C and CIFAR100C, VCoTTA achieves consistent and clear improvements over baselines. For example, VCoTTA reduces the error rate to 13.1% compared with 15.3% from SWA on CIFAR10C, and to 28.4% compared with 30.0% from SWA on CIFAR100C. These gains highlight the effectiveness of explicitly modeling uncertainty and leveraging the prior mixture mechanism, particularly in settings where long-term adaptation and moderately large label spaces make error accumulation a critical challenge. On the more challenging ImageNetC benchmark with 1,000 classes (Fig. 3), VCoTTA delivers competitive results, obtaining 64.2% compared with 66.7% from CoTTA. A complementary stacked analysis of corruption groups further shows that VCoTTA distributes errors more evenly, suggesting that it accumulates errors more slowly under diverse and long-term shifts. These results confirm that priors inevitably drift during continual adaptation. By explicitly modeling uncertainty and adaptively combining priors, VCoTTA slows down this drift and improves robustness. The performance gains are especially pronounced on CIFAR10C and CIFAR100C, which serve as strong evidence of the method’s effectiveness.
VI-C Ablation Study on Auxiliary Components
To better understand the contribution of the auxiliary components, we conduct an ablation study while keeping the main framework of VCoTTA fixed. The backbone always includes variational inference with the uncertainty-aware prior mixture. We specifically examine the effect of the Variational Warm-Up (VWU) and the Symmetric Cross-Entropy (SCE). In Table II, a checkmark indicates that the component is included, while absence means it is removed. In particular, “w/o VWU” means the BNN is trained directly without the warm-up stage, and “w/o SCE” means the symmetric cross-entropy is replaced with the standard cross-entropy loss. The results show that both components provide additional improvements on top of the core framework. Removing VWU causes a clear drop in performance. For example, CIFAR10C decreases from 13.9% to 18.4%, and CIFAR100C decreases from 28.8% to 31.5%. This demonstrates that stochastic dynamics introduced during warm-up are important for stable adaptation. Adding SCE further improves accuracy on CIFAR10C and CIFAR100C, because it balances gradients between high-confidence and low-confidence predictions. On ImageNetC, the effect of SCE is marginal, which may be related to class imbalance and sensitivity in large-scale data. These findings confirm that VWU and SCE are effective auxiliary enhancements, while the main contribution of VCoTTA remains in the variational Bayesian formulation and the adaptive prior mixture.
Figure 3: Sum of error rate (%) on ImageNet-to-ImageNetC.
TABLE II: Ablation study of auxiliarxny components under corruption severity level 5. All variants keep the core framework of variational inference with the uncertainty-aware prior mixture unchanged. A checkmark indicates that the auxiliary component is used, while absence means it is removed.
VCoTTA VWU SCE CIFAR10C CIFAR100C ImageNetC
$\surd$ 18.4 31.5 68.1
$\surd$ $\surd$ 17.1 31.2 68.3
$\surd$ $\surd$ 13.9 28.8 64.2
$\surd$ $\surd$ $\surd$ 13.1 28.4 64.7
TABLE III: Different weights for mixture of priors.
No. $\alpha$ $1-\alpha$ CIFAR10C CIFAR100C ImageNetC
@@ VI-C Ablation Study on Auxiliary Components
4 Eq. (11) 13.1 28.4 64.7
VI-D Mixture of Priors
In Sec. IV-B, we introduce a Gaussian mixture strategy, where the current prior is approximated as the weighted sum of the source and teacher priors. The weights are determined by computing the entropy over multiple augmentations of two models. To assess the effectiveness of these weights, we compare them with three naive weighting configurations: using only the source model, using only the teacher model, and a simple average with equal weights for both models. The results, shown in Table III, reveal that relying solely on the source model or the teacher model (i.e., weighting with $(1,0)$ and $(0,1)$) results in suboptimal performance. Additionally, naive weighting with equal contributions from both models (i.e., $(0.5,0.5)$) proves ineffective for CTTA due to the inherent uncertainty in both models. In contrast, the proposed adaptive weights for the Gaussian mixture demonstrate its effectiveness. This underscores the significance of striking a balance between the two prior models in an unsupervised environment. The trade-off implies the need to discern when the source model’s knowledge is more applicable and when the teacher model’s shifting knowledge takes precedence.
VI-E Uncertainty Estimation
To evaluate the uncertainty estimation, we use negative loglikelihood (NLL) and Brier Score (BS) [brier1950verification]. Both NLL and BS are proper scoring rules [gneiting2007strictly], and they are minimized if and only if the predicted distribution becomes identical to the actual distribution:
@@ VI-E Uncertainty Estimation
where $\mathcal{D}^{\text{test}}$ denotes the test set, i.e., the unsupervised test dataset $\mathcal{U}$ with labels. We evaluate NLL and BS with a severity level of 5 for all corruption types, and the compared results with SOTAs are shown in Table IV. We have the following observations. First, most methods suffer from low confidence in terms of NLL and BS because of the drift priors, where the model is unreliable gradually, and the error accumulation makes the model perform poorly. Our approach outperforms most other approaches in terms of NLL and BS, demonstrating the superiority in improving uncertainty estimation. We also find that PETAL [brahma2023probabilistic] shows good NLL and BS, because PETAL forces the prediction over-confident to unreliable priors, thus PETAL shows unsatisfactory results on adaptation accuracy, such as 31.5% vs. 28.4% (Ours) on CIFAR100C.
TABLE IV: Uncertainty estimation via NLL and BS.
@@
TABLE V: Gradually changing on severity 5.
Method CIFAR10C CIFAR100C ImageNetC
@@
VI-F Gradually Corruption
We also show gradual corruption results instead of constant severity in the major comparison, and the results are reported in Table V. Specifically, each corruption adopts the gradual changing sequence: $1\rightarrow 2\rightarrow 3\rightarrow 4\rightarrow 5\rightarrow 4\rightarrow 3\rightarrow 2\rightarrow 1$, where the severity level is the lowest 1 when corruption type changes, therefore, the type change is gradual. The distribution shift within each type is also gradual. Under this situation, our VCoTTA also outperforms other methods, such as 8.9% vs. 10.5% (PETAL) on CIFAR10C, and 24.4% vs. 26.3% (CoTTA) on CIFAR100C. The results show that the proposed VCoTTA based on BNN is also effective when the distribution change is uncertain.
VI-G Augmentation Analysis
In our method, we use the standard augmentation following CoTTA [wang2022continual]. In this subsection, we analyze the some characteristics via experiments.
VI-G1 Confidence margin in augmentations
First, we analyze the margin $\epsilon$ in Eq. (14). We experimentally validate different margins with more choices. Experimental results are shown in Tables VI. The results indicate that different datasets may require different margins to control confidence. Moreover, Eq. (14) signifies that the reliable teacher likelihood is represented by the mean of its augmentations with $\epsilon$ more confidence than the teacher itself. Tables VI illustrates the selection of $\epsilon$ in our approach on CIFAR10C, CIFAR100C and ImageNetC. Note that when $\epsilon=-1$, it means no margin is used and the method will use all augmentated samples, i.e., without using Eq. (14). The results show that the proposed margin can effectively filter out unreliable augmented samples and achieve a better teacher log-likelihood.
TABLE VI: Analysis on confidence margin.